# Predicting perturbation targets with causal differential networks

### v1

@@ 20 Ut-Igsp 1 78 19 26
10 10 Ut-Igsp $0.29\scriptstyle\pm.02$ $0.57\scriptstyle\pm.03$ $0.30\scriptstyle\pm.01$ $0.63\scriptstyle\pm.02$ 0.33±.04plus-or-minus0.33.040.$3$04 $0.59\scriptstyle\pm.03$ $0.32\scriptstyle\pm.01$ $0.64\scriptstyle\pm.02$
Dcdi-G $0.39\scriptstyle\pm.04$ $0.52\scriptstyle\pm.04$ $40$04 $0.51\scriptstyle\pm.04$ $0.38\scriptstyle\pm.03$ $0.49\scriptstyle\pm.02$ $0.36\scriptstyle\pm.03$ $0.49\scriptstyle\pm.03$
Dcdi-Dsf $0.36\scriptstyle\pm.02$ $0.51\scriptstyle\pm.04$ $0.40\scriptstyle\pm.05$ $0.53\scriptstyle\pm.05$ $0.37\scriptstyle\pm.02$ $0.50\scriptstyle\pm.01$ $0.38\scriptstyle\pm.02$ $0.49\scriptstyle\pm.03$
Dci 0.55±.08plus-or-minus0.55.080.$5$08 $0.79\scriptstyle\pm.04$ $\textbf{0.50}\scriptstyle\pm.08$ $\textbf{0.70}\scriptstyle\pm.05$ $0.47\scriptstyle\pm.09$ $0.71\scriptstyle\pm.06$ $0.41\scriptstyle\pm.02$ $0.67\scriptstyle\pm.01$
BaCaDI-E $0.26\scriptstyle\pm.03$ $0.61\scriptstyle\pm.05$ $0.42\scriptstyle\pm.13$ $0.63\scriptstyle\pm.11$ $0.25\scriptstyle\pm.03$ $0.61\scriptstyle\pm.05$ $0.30\scriptstyle\pm.05$ $0.61\scriptstyle\pm.04$
BaCaDI-M $0.23\scriptstyle\pm.01$ $0.56\scriptstyle\pm.03$ $0.37\scriptstyle\pm.11$ $0.62\scriptstyle\pm.11$ 0.22±.02plus-or-minus0.22.020.$2$02 $0.56\scriptstyle\pm.04$ $0.30\scriptstyle\pm.05$ $0.61\scriptstyle\pm.04$
[…]
Cdn (Axial) $\textbf{0.73}\scriptstyle\pm.05$ $\textbf{0.85}\scriptstyle\pm.03$ $0.48\scriptstyle\pm.06$ $0.64\scriptstyle\pm.05$ $\textbf{0.62}\scriptstyle\pm.09$ 0.77±.07plus-or-minus0.77.070.$7$07 $\textbf{0.73}\scriptstyle\pm.07$ $\textbf{0.82}\scriptstyle\pm.06$
10 20 Ut-Igsp $0.28\scriptstyle\pm.02$ $0.56\scriptstyle\pm.03$ $0.26\scriptstyle\pm.01$ $0.59\scriptstyle\pm.03$ $0.28\scriptstyle\pm.03$ $0.58\scriptstyle\pm.04$ $0.27\scriptstyle\pm.01$ $0.60\scriptstyle\pm.01$
Dcdi-G $0.42\scriptstyle\pm.02$ $0.54\scriptstyle\pm.02$ $0.40\scriptstyle\pm.03$ $0.51\scriptstyle\pm.04$ $0.40\scriptstyle\pm.06$ $0.53\scriptstyle\pm.05$ $0.39\scriptstyle\pm.02$ $0.51\scriptstyle\pm.03$
Dcdi-Dsf $0.41\scriptstyle\pm.03$ $0.52\scriptstyle\pm.03$ $0.39\scriptstyle\pm.05$ $0.52\scriptstyle\pm.05$ $0.39\scriptstyle\pm.02$ $0.51\scriptstyle\pm.02$ $0.39\scriptstyle\pm.03$ $50$05
[…]
BaCaDI-E $0.34\scriptstyle\pm.04$ $0.68\scriptstyle\pm.03$ $0.53\scriptstyle\pm.08$ $0.71\scriptstyle\pm.04$ 0.33±.03plus-or-minus0.33.030.$3$03 $0.72\scriptstyle\pm.04$ $0.64\scriptstyle\pm.07$ $0.78\scriptstyle\pm.05$
BaCaDI-M $0.27\scriptstyle\pm.02$ $0.63\scriptstyle\pm.03$ $0.48\scriptstyle\pm.09$ $0.71\scriptstyle\pm.04$ $0.27\scriptstyle\pm.04$ $0.64\scriptstyle\pm.08$ $0.59\scriptstyle\pm.09$ 0.77±.06plus-or-minus0.77.060.$7$06
[…]
20 20 Ut-Igsp $0.15\scriptstyle\pm.01$ $0.54\scriptstyle\pm.01$ $0.18\scriptstyle\pm.01$ $0.65\scriptstyle\pm.02$ $20$02 $0.60\scriptstyle\pm.02$ $0.21\scriptstyle\pm.02$ $0.67\scriptstyle\pm.02$
Dcdi-G $0.24\scriptstyle\pm.02$ $0.49\scriptstyle\pm.02$ 0.22±.02plus-or-minus0.22.020.$2$02 $0.49\scriptstyle\pm.02$ 0.22±.03plus-or-minus0.22.030.$2$03 $0.50\scriptstyle\pm.03$ 0.22±.02plus-or-minus0.22.020.$2$02 $0.49\scriptstyle\pm.03$
Dcdi-Dsf $0.24\scriptstyle\pm.02$ $0.50\scriptstyle\pm.02$ $0.23\scriptstyle\pm.02$ $0.50\scriptstyle\pm.03$ 0.22±.01plus-or-minus0.22.010.$2$01 $0.50\scriptstyle\pm.02$ $0.20\scriptstyle\pm.01$ $0.48\scriptstyle\pm.02$
Dci $0.43\scriptstyle\pm.03$ 0.77±.01plus-or-minus0.77.010.$7$01 $0.45\scriptstyle\pm.08$ $\textbf{0.75}\scriptstyle\pm.04$ $0.48\scriptstyle\pm.05$ $0.76\scriptstyle\pm.03$ $0.50\scriptstyle\pm.02$ $0.79\scriptstyle\pm.02$
BaCaDI-E $0.14\scriptstyle\pm.01$ $0.61\scriptstyle\pm.03$ $0.25\scriptstyle\pm.10$ $60$06 0.22±.06plus-or-minus0.22.060.$2$06 $0.76\scriptstyle\pm.06$ $0.30\scriptstyle\pm.06$ $0.74\scriptstyle\pm.05$
BaCaDI-M $0.12\scriptstyle\pm.00$ $0.58\scriptstyle\pm.02$ $0.17\scriptstyle\pm.03$ $0.60\scriptstyle\pm.05$ $0.14\scriptstyle\pm.02$ 0.66±.05plus-or-minus0.66.050.$6$05 0.22±.05plus-or-minus0.22.050.$2$05 $0.71\scriptstyle\pm.06$
Cdn (Mlp) $\textbf{0.67}\scriptstyle\pm.05$ $\textbf{0.87}\scriptstyle\pm.02$ $0.28\scriptstyle\pm.01$ $0.54\scriptstyle\pm.03$ $60$06 $\textbf{0.84}\scriptstyle\pm.03$ $0.34\scriptstyle\pm.04$ $0.51\scriptstyle\pm.04$
Cdn (Axial) $0.65\scriptstyle\pm.03$ $0.86\scriptstyle\pm.01$ $\textbf{0.49}\scriptstyle\pm.05$ $0.72\scriptstyle\pm.03$ $0.58\scriptstyle\pm.05$ $0.83\scriptstyle\pm.02$ $\textbf{0.65}\scriptstyle\pm.02$ $\textbf{0.81}\scriptstyle\pm.02$
20 40 Ut-Igsp $0.14\scriptstyle\pm.01$ $0.57\scriptstyle\pm.01$ $0.14\scriptstyle\pm.01$ $0.61\scriptstyle\pm.02$ $0.20\scriptstyle\pm.01$ $0.62\scriptstyle\pm.02$ $0.19\scriptstyle\pm.01$ $0.65\scriptstyle\pm.01$
Dcdi-G 0.22±.03plus-or-minus0.22.030.$2$03 $0.49\scriptstyle\pm.03$ $0.21\scriptstyle\pm.02$ $0.48\scriptstyle\pm.03$ 0.22±.02plus-or-minus0.22.020.$2$02 $0.49\scriptstyle\pm.02$ 0.22±.02plus-or-minus0.22.020.$2$02 $0.50\scriptstyle\pm.02$
Dcdi-Dsf $0.21\scriptstyle\pm.02$ $0.46\scriptstyle\pm.03$ $0.21\scriptstyle\pm.02$ $0.47\scriptstyle\pm.02$ $0.23\scriptstyle\pm.02$ $0.50\scriptstyle\pm.03$ $0.21\scriptstyle\pm.01$ $0.49\scriptstyle\pm.03$
[…]
BaCaDI-E $0.23\scriptstyle\pm.05$ $0.70\scriptstyle\pm.04$ 0.33±.10plus-or-minus0.33.100.$3$10 $0.67\scriptstyle\pm.04$ $0.29\scriptstyle\pm.08$ $0.83\scriptstyle\pm.04$ $0.45\scriptstyle\pm.18$ $0.78\scriptstyle\pm.06$
BaCaDI-M $0.15\scriptstyle\pm.01$ $0.64\scriptstyle\pm.02$ $0.19\scriptstyle\pm.04$ $0.63\scriptstyle\pm.02$ $0.17\scriptstyle\pm.02$ $0.70\scriptstyle\pm.05$ $0.31\scriptstyle\pm.10$ 0.77±.07plus-or-minus0.77.070.$7$07
[…]
Cdn (Axial) $0.74\scriptstyle\pm.05$ $0.90\scriptstyle\pm.02$ $0.38\scriptstyle\pm.03$ $0.65\scriptstyle\pm.01$ $0.68\scriptstyle\pm.08$ $0.87\scriptstyle\pm.04$ $\textbf{0.64}\scriptstyle\pm.03$ $0.79\scriptstyle\pm.03$

### v2

@@ 20 Ut-Igsp 1 78 19 26
10 10 Ut-Igsp .$29\scriptstyle\pm.02$$57\scriptstyle\pm.03$$30\scriptstyle\pm.01$$63\scriptstyle\pm.02$$3$04 .$59\scriptstyle\pm.03$$32\scriptstyle\pm.01$$64\scriptstyle\pm.02$$28\scriptstyle\pm.03$$58\scriptstyle\pm.04$$26\scriptstyle\pm.02$$59\scriptstyle\pm.03$
Dcdi-G .$39\scriptstyle\pm.04$$52\scriptstyle\pm.04$$40\scriptstyle\pm.04$$51\scriptstyle\pm.04$$38\scriptstyle\pm.03$$49\scriptstyle\pm.02$$36\scriptstyle\pm.03$$49\scriptstyle\pm.03$$37\scriptstyle\pm.03$$49\scriptstyle\pm.03$$40\scriptstyle\pm.04$$52\scriptstyle\pm.06$
Dcdi-Dsf .$36\scriptstyle\pm.02$$51\scriptstyle\pm.04$$40\scriptstyle\pm.05$$53\scriptstyle\pm.05$$37\scriptstyle\pm.02$$50\scriptstyle\pm.01$$38\scriptstyle\pm.02$$49\scriptstyle\pm.03$$37\scriptstyle\pm.02$$50\scriptstyle\pm.02$$37\scriptstyle\pm.02$$48\scriptstyle\pm.03$
Bacadi-E .$26\scriptstyle\pm.03$$61\scriptstyle\pm.05$$42\scriptstyle\pm.13$$63\scriptstyle\pm.11$$25\scriptstyle\pm.03$$61\scriptstyle\pm.05$$30\scriptstyle\pm.05$$61\scriptstyle\pm.04$$28\scriptstyle\pm.04$$64\scriptstyle\pm.06$$43\scriptstyle\pm.12$$69\scriptstyle\pm.09$
Bacadi-M .$23\scriptstyle\pm.01$$56\scriptstyle\pm.03$$37\scriptstyle\pm.11$$62\scriptstyle\pm.11$$2$02 .$56\scriptstyle\pm.04$$30\scriptstyle\pm.05$$61\scriptstyle\pm.04$$25\scriptstyle\pm.03$$61\scriptstyle\pm.05$$42\scriptstyle\pm.11$$69\scriptstyle\pm.09$
Dci .55±.08plus-or-minus.55.08.$5$08 .$79\scriptstyle\pm.04$$50\scriptstyle\pm.08$$70\scriptstyle\pm.05$$47\scriptstyle\pm.09$$71\scriptstyle\pm.06$$41\scriptstyle\pm.02$$67\scriptstyle\pm.01$$59\scriptstyle\pm.07$$80\scriptstyle\pm.05$$51\scriptstyle\pm.05$$76\scriptstyle\pm.04$
Mb+Ci .$53\scriptstyle\pm.05$$64\scriptstyle\pm.03$$57\scriptstyle\pm.10$$6$06 .$70\scriptstyle\pm.09$$79\scriptstyle\pm.07$$85\scriptstyle\pm.07$$89\scriptstyle\pm.05$$84\scriptstyle\pm.08$$8$06 .$96\scriptstyle\pm.05$$97\scriptstyle\pm.04$
Cdn .$86\scriptstyle\pm.02$$93\scriptstyle\pm.01$$74\scriptstyle\pm.06$$85\scriptstyle\pm.06$$83\scriptstyle\pm.09$$91\scriptstyle\pm.05$$89\scriptstyle\pm.05$$94\scriptstyle\pm.03$$8$06 .$94\scriptstyle\pm.04$$85\scriptstyle\pm.07$$91\scriptstyle\pm.05$
Cdn (no $\mathcal{L}_{G}$) .$67\scriptstyle\pm.09$$78\scriptstyle\pm.07$$82\scriptstyle\pm.05$$89\scriptstyle\pm.05$$82\scriptstyle\pm.08$$91\scriptstyle\pm.05$$91\scriptstyle\pm.04$$96\scriptstyle\pm.02$$89\scriptstyle\pm.05$$95\scriptstyle\pm.03$$94\scriptstyle\pm.03$$97\scriptstyle\pm.02$
Cdn (full) .$87\scriptstyle\pm.01$$94\scriptstyle\pm.01$$9$01 $1.0\scriptstyle\pm.00$$82\scriptstyle\pm.08$$91\scriptstyle\pm.05$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$$95\scriptstyle\pm.05$$98\scriptstyle\pm.02$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$
10 20 Ut-Igsp .$28\scriptstyle\pm.02$$56\scriptstyle\pm.03$$26\scriptstyle\pm.01$$59\scriptstyle\pm.03$$28\scriptstyle\pm.03$$58\scriptstyle\pm.04$$27\scriptstyle\pm.01$$60\scriptstyle\pm.01$$25\scriptstyle\pm.00$$5$03 .$25\scriptstyle\pm.02$$56\scriptstyle\pm.05$
Dcdi-G .$42\scriptstyle\pm.02$$54\scriptstyle\pm.02$$40\scriptstyle\pm.03$$51\scriptstyle\pm.04$$40\scriptstyle\pm.06$$53\scriptstyle\pm.05$$39\scriptstyle\pm.02$$51\scriptstyle\pm.03$$37\scriptstyle\pm.02$$50\scriptstyle\pm.01$$35\scriptstyle\pm.03$$49\scriptstyle\pm.04$
Dcdi-Dsf .$41\scriptstyle\pm.03$$52\scriptstyle\pm.03$$39\scriptstyle\pm.05$$52\scriptstyle\pm.05$$39\scriptstyle\pm.02$$51\scriptstyle\pm.02$$39\scriptstyle\pm.03$$50\scriptstyle\pm.05$$36\scriptstyle\pm.03$$50\scriptstyle\pm.05$$36\scriptstyle\pm.03$$49\scriptstyle\pm.03$
Bacadi-E .$34\scriptstyle\pm.04$$68\scriptstyle\pm.03$$53\scriptstyle\pm.08$$71\scriptstyle\pm.04$$3$03 .$72\scriptstyle\pm.04$$64\scriptstyle\pm.07$$78\scriptstyle\pm.05$$34\scriptstyle\pm.05$$71\scriptstyle\pm.04$$51\scriptstyle\pm.07$$73\scriptstyle\pm.04$
Bacadi-M .$27\scriptstyle\pm.02$$63\scriptstyle\pm.03$$48\scriptstyle\pm.09$$71\scriptstyle\pm.04$$27\scriptstyle\pm.04$$64\scriptstyle\pm.08$$59\scriptstyle\pm.09$$7$06 .$28\scriptstyle\pm.03$$65\scriptstyle\pm.05$$45\scriptstyle\pm.08$$71\scriptstyle\pm.05$
Dci .$59\scriptstyle\pm.03$$78\scriptstyle\pm.03$$57\scriptstyle\pm.04$$7$03 .$68\scriptstyle\pm.07$$82\scriptstyle\pm.04$$65\scriptstyle\pm.08$$84\scriptstyle\pm.05$$70\scriptstyle\pm.03$$84\scriptstyle\pm.02$$6$06 .$83\scriptstyle\pm.05$
Mb+Ci .$54\scriptstyle\pm.11$$6$09 .$49\scriptstyle\pm.06$$62\scriptstyle\pm.04$$83\scriptstyle\pm.04$$87\scriptstyle\pm.03$$96\scriptstyle\pm.02$$96\scriptstyle\pm.02$$80\scriptstyle\pm.09$$84\scriptstyle\pm.08$$96\scriptstyle\pm.02$$96\scriptstyle\pm.02$
Cdn .88±.04plus-or-minus.88.04.$8$04 .$95\scriptstyle\pm.02$$56\scriptstyle\pm.07$$71\scriptstyle\pm.07$$96\scriptstyle\pm.02$$98\scriptstyle\pm.01$$95\scriptstyle\pm.04$$97\scriptstyle\pm.02$$94\scriptstyle\pm.04$$98\scriptstyle\pm.02$$68\scriptstyle\pm.02$$79\scriptstyle\pm.02$
Cdn (no $\mathcal{L}_{G}$) .$79\scriptstyle\pm.06$$90\scriptstyle\pm.04$$7$04 .$86\scriptstyle\pm.02$$94\scriptstyle\pm.01$$97\scriptstyle\pm.01$$95\scriptstyle\pm.05$$96\scriptstyle\pm.04$$87\scriptstyle\pm.02$$93\scriptstyle\pm.02$$94\scriptstyle\pm.02$$96\scriptstyle\pm.01$
Cdn (full) .$89\scriptstyle\pm.03$$94\scriptstyle\pm.02$$98\scriptstyle\pm.02$$9$01 .$97\scriptstyle\pm.01$$9$00 $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$$96\scriptstyle\pm.03$$98\scriptstyle\pm.01$$98\scriptstyle\pm.02$$9$01
20 20 Ut-Igsp .$15\scriptstyle\pm.01$$54\scriptstyle\pm.01$$18\scriptstyle\pm.01$$65\scriptstyle\pm.02$$20\scriptstyle\pm.02$$60\scriptstyle\pm.02$$21\scriptstyle\pm.02$$67\scriptstyle\pm.02$$15\scriptstyle\pm.01$$57\scriptstyle\pm.02$$17\scriptstyle\pm.01$$65\scriptstyle\pm.01$
Dcdi-G .$24\scriptstyle\pm.02$$49\scriptstyle\pm.02$$2$02 .$49\scriptstyle\pm.02$$2$03 .$50\scriptstyle\pm.03$$2$02 .$49\scriptstyle\pm.03$$2$02 .$50\scriptstyle\pm.04$$2$02 .$50\scriptstyle\pm.03$
Dcdi-Dsf .$24\scriptstyle\pm.02$$50\scriptstyle\pm.02$$23\scriptstyle\pm.02$$50\scriptstyle\pm.03$$2$01 .$50\scriptstyle\pm.02$$20\scriptstyle\pm.01$$48\scriptstyle\pm.02$$2$02 .$50\scriptstyle\pm.04$$21\scriptstyle\pm.02$$47\scriptstyle\pm.03$
Bacadi-E .$14\scriptstyle\pm.01$$61\scriptstyle\pm.03$$25\scriptstyle\pm.10$$60\scriptstyle\pm.06$$2$06 .$76\scriptstyle\pm.06$$30\scriptstyle\pm.06$$74\scriptstyle\pm.05$$15\scriptstyle\pm.02$$65\scriptstyle\pm.03$$17\scriptstyle\pm.05$$57\scriptstyle\pm.05$
Bacadi-M .$12\scriptstyle\pm.00$$58\scriptstyle\pm.02$$17\scriptstyle\pm.03$$60\scriptstyle\pm.05$$14\scriptstyle\pm.02$$6$05 .22±.05plus-or-minus.22.05.$2$05 .$71\scriptstyle\pm.06$$12\scriptstyle\pm.01$$59\scriptstyle\pm.04$$14\scriptstyle\pm.02$$56\scriptstyle\pm.05$
Dci .$43\scriptstyle\pm.03$$7$01 .$45\scriptstyle\pm.08$$75\scriptstyle\pm.04$$48\scriptstyle\pm.05$$76\scriptstyle\pm.03$$50\scriptstyle\pm.02$$79\scriptstyle\pm.02$$46\scriptstyle\pm.05$$75\scriptstyle\pm.02$$38\scriptstyle\pm.12$$73\scriptstyle\pm.06$
Mb+Ci .$48\scriptstyle\pm.09$$6$05 .$57\scriptstyle\pm.08$$70\scriptstyle\pm.05$$72\scriptstyle\pm.10$$84\scriptstyle\pm.06$$92\scriptstyle\pm.04$$94\scriptstyle\pm.02$$68\scriptstyle\pm.06$$80\scriptstyle\pm.04$$89\scriptstyle\pm.03$$92\scriptstyle\pm.02$
Cdn .$74\scriptstyle\pm.02$$90\scriptstyle\pm.01$$85\scriptstyle\pm.06$$96\scriptstyle\pm.02$$78\scriptstyle\pm.09$$92\scriptstyle\pm.04$$92\scriptstyle\pm.02$$98\scriptstyle\pm.00$$75\scriptstyle\pm.06$$91\scriptstyle\pm.03$$78\scriptstyle\pm.06$$92\scriptstyle\pm.04$
Cdn (no $\mathcal{L}_{G}$) .$50\scriptstyle\pm.08$$73\scriptstyle\pm.07$$87\scriptstyle\pm.03$$96\scriptstyle\pm.01$$76\scriptstyle\pm.09$$90\scriptstyle\pm.05$$92\scriptstyle\pm.03$$98\scriptstyle\pm.01$$71\scriptstyle\pm.04$$89\scriptstyle\pm.02$$93\scriptstyle\pm.02$$9$00
Cdn (full) .$75\scriptstyle\pm.03$$91\scriptstyle\pm.02$$9$01 $1.0\scriptstyle\pm.00$$79\scriptstyle\pm.08$$93\scriptstyle\pm.04$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$$76\scriptstyle\pm.05$$92\scriptstyle\pm.02$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$
20 40 Ut-Igsp .$14\scriptstyle\pm.01$$57\scriptstyle\pm.01$$14\scriptstyle\pm.01$$61\scriptstyle\pm.02$$20\scriptstyle\pm.01$$62\scriptstyle\pm.02$$19\scriptstyle\pm.01$$65\scriptstyle\pm.01$$13\scriptstyle\pm.01$$56\scriptstyle\pm.02$$14\scriptstyle\pm.00$$60\scriptstyle\pm.01$
Dcdi-G .22±.03plus-or-minus.22.03.$2$03 .$49\scriptstyle\pm.03$$21\scriptstyle\pm.02$$48\scriptstyle\pm.03$$2$02 .$49\scriptstyle\pm.02$$2$02 .$50\scriptstyle\pm.02$$21\scriptstyle\pm.01$$48\scriptstyle\pm.02$$23\scriptstyle\pm.01$$51\scriptstyle\pm.02$
Dcdi-Dsf .$21\scriptstyle\pm.02$$46\scriptstyle\pm.03$$21\scriptstyle\pm.02$$47\scriptstyle\pm.02$$23\scriptstyle\pm.02$$50\scriptstyle\pm.03$$21\scriptstyle\pm.01$$49\scriptstyle\pm.03$$21\scriptstyle\pm.02$$49\scriptstyle\pm.02$$21\scriptstyle\pm.03$$49\scriptstyle\pm.02$
Bacadi-E .$23\scriptstyle\pm.05$$70\scriptstyle\pm.04$$3$10 .$67\scriptstyle\pm.04$$29\scriptstyle\pm.08$$83\scriptstyle\pm.04$$45\scriptstyle\pm.18$$78\scriptstyle\pm.06$$20\scriptstyle\pm.05$$73\scriptstyle\pm.05$$51\scriptstyle\pm.13$$79\scriptstyle\pm.04$
Bacadi-M .$15\scriptstyle\pm.01$$64\scriptstyle\pm.02$$19\scriptstyle\pm.04$$63\scriptstyle\pm.02$$17\scriptstyle\pm.02$$70\scriptstyle\pm.05$$31\scriptstyle\pm.10$$7$07 .$14\scriptstyle\pm.01$$65\scriptstyle\pm.03$$32\scriptstyle\pm.06$$7$04
Dci .$60\scriptstyle\pm.06$$85\scriptstyle\pm.03$$50\scriptstyle\pm.02$$7$03 .$56\scriptstyle\pm.08$$81\scriptstyle\pm.05$$59\scriptstyle\pm.06$$84\scriptstyle\pm.04$$64\scriptstyle\pm.06$$87\scriptstyle\pm.02$$6$05 .88±.03plus-or-minus.88.03.$8$03
Mb+Ci .$53\scriptstyle\pm.09$$70\scriptstyle\pm.06$$47\scriptstyle\pm.06$$64\scriptstyle\pm.03$$78\scriptstyle\pm.04$$8$02 .$95\scriptstyle\pm.03$$96\scriptstyle\pm.02$$83\scriptstyle\pm.03$$89\scriptstyle\pm.02$$96\scriptstyle\pm.04$$97\scriptstyle\pm.03$
Cdn .$86\scriptstyle\pm.04$$96\scriptstyle\pm.02$$79\scriptstyle\pm.09$$93\scriptstyle\pm.05$$8$03 .$97\scriptstyle\pm.01$$96\scriptstyle\pm.02$$9$00 .88±.03plus-or-minus.88.03.$8$03 .$97\scriptstyle\pm.01$$71\scriptstyle\pm.05$$87\scriptstyle\pm.04$
Cdn (no $\mathcal{L}_{G}$) .$69\scriptstyle\pm.09$$85\scriptstyle\pm.07$$80\scriptstyle\pm.06$$94\scriptstyle\pm.03$$84\scriptstyle\pm.03$$94\scriptstyle\pm.01$$95\scriptstyle\pm.02$$9$00 .$84\scriptstyle\pm.05$$93\scriptstyle\pm.03$$91\scriptstyle\pm.02$$98\scriptstyle\pm.00$
Cdn (full) .$89\scriptstyle\pm.06$$97\scriptstyle\pm.02$$9$01 $1.0\scriptstyle\pm.00$$89\scriptstyle\pm.03$$97\scriptstyle\pm.01$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$$93\scriptstyle\pm.03$$98\scriptstyle\pm.01$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$